### v1

# From Grounding to Manipulation: Case Studies of Foundation Model Integration in Embodied Robotic Systems

Abstract
Foundation models (FMs) are increasingly used to bridge language and action in embodied agents, yet the operational characteristics of different FM integration strategies remain under-explored—particularly for complex instruction following and versatile action generation in changing environments. This paper examines three paradigms for building robotic systems: end-to-end vision-language-action (VLA) models that implicitly integrate perception and planning, and modular pipelines incorporating either vision-language models (VLMs) or multimodal large language models (LLMs). We evaluate these paradigms through two focused case studies: a complex instruction grounding task assessing fine-grained instruction understanding and cross-modal disambiguation, and an object manipulation task targeting skill transfer via VLA finetuning. Our experiments in zero-shot and few-shot settings reveal trade-offs in generalization and data efficiency. By exploring performance limits, we distill design implications for developing language-driven physical agents and outline emerging challenges and opportunities for FM-powered robotics in real-world conditions.
From Grounding to Manipulation: Case Studies of Foundation Model Integration in Embodied Robotic Systems
Xiuchao Sui1thanks: xiuchao.sui@gmail.com.  and Daiying Tian1thanks: Equal contribution. Corresponding author.  and Qi Sun 2  and Ruirui Chen1  and Dongkyu Choi 1and Kenneth Kwok 1and Soujanya Poria 2 1Institute of High Performance Computing, A*STAR, Singapore 2Singapore University of Technology and Design, Singapore
1 Introduction
Natural language is emerging as a universal interface for embodied robotic systems. Recent foundation models (FMs) allow robots to follow free-form instructions in perception, reasoning, and motor commands, offering the promise of language-grounded autonomy. They include multimodal large language models (LLMs) Grattafiori et al. (2024); Bai et al. (2025); Lu et al. (2024), vision–language models (VLMs) Liu et al. (2024a); Ravi et al. (2025); Ren et al. (2024); Li et al. (2023), and vision–language–action (VLA) models Kim et al. (2024); Zheng et al. (2025); Qu et al. (2025); Bu et al. (2025).
However, turning the promise of language-grounded autonomy into deployable systems is highly challenging. Robots must (i) map ambiguous instructions to the physical world (instruction grounding), (ii) execute reliably across novel objects, scenes, and robot morphologies (generalizable execution), and (iii) achieving the aforementioned goals with limited data (efficient adaptation). How well different FM integration strategies meet these competing requirements remains under-explored (Fig. 1).
Figure 1: Challenges of foundation models in embodied robotic systems include cross-modal instruction grounding, generalization across environments and morphologies, and data-efficient adaptation for the real world.
To our best knowledge, this work delivers the first head-to-head empirical comparison of three prevalent integration paradigms: end-to-end VLAs that directly map language and vision to actions, multimodal LLM agents that orchestrate perception and control through tool calls, and modular VLM pipelines that couple perception-specialist FMs with task-specific planners (Fig. 2; Table 1).
We evaluate these paradigms through tabletop case studies designed to highlight their complementary strengths and limitations. Specifically, we consider two task categories: (i) Complex Instruction Grounding, which probes fine-grained understanding and cross-modal disambiguation (Sec.3); and (ii) Object Manipulation, which measures the ability to transfer learned skills after VLA fine-tuning under distribution shifts, complemented by comparative and ablation studies (Sec.4).
Our zero-shot grounding experiments reveal distinct trade-offs across integration strategies. VLM pipelines prioritize interpretability and data efficiency, sacrificing flexibility and peak performance. While underperforming on handling complex instruction grounding, they deliver moderate performance in object grounding—using less than 1% of the parameters required by multimodal LLMs. In contrast, multimodal LLM agents generalize better on complex instructions but incur significantly higher inference costs. Notably, smaller reasoning-focused models such as GPT-4o-mini can even outperform larger models like GPT-4o on certain tasks. VLAs, with their tightly coupled perception-to-action pathways, support streamlined action generation, yet struggle to reason about rare or abstract concepts. We further examine the trade-off between model size and performance by analyzing quantization effects on open-source multimodal LLMs. These findings offer practical guidance for developing language-driven robotic systems under real-world constraints.
Within the VLA paradigm, we categorize models by their action generation mechanisms—autoregressive Vaswani et al. (2017); Kim et al. (2024); Hung et al. (2025) and diffusion-based approaches Ho et al. (2020); Chi et al. (2023); Reuss et al. (2024); Wen et al. (2024). We evaluate their adaptability through fine-tuning under distribution shifts, mirroring real-world deployment scenarios. To assess generalization, we analyze robustness to environmental perturbations and variation in object appearance and robot morphology.
To summarize, the main contributions of this work are as follows:
  • To our best knowledge, we present the first systematic comparison of end-to-end VLA, modular VLM, and modular multimodal LLM architectures on the same set of embodied tasks.
  • We release a dataset and accompanying code [^1] https://github.com/xiuchao/InstructionGrounding that supports the evaluation of complex instruction grounding and manipulation transfer—covering accordance comprehension, cross-modal reasoning, and motor adaptation.
  • We benchmark state-of-the-art VLAs and multimodal LLMs, offering timely insight into the capabilities and failure modes of modern FMs in robotics.
  • We distill actionable trade-offs that practitioners can apply when choosing an FM stack for language-driven embodied agents.
  • We release a complete, end-to-end claw-machine robot system to demonstrate FM integrations in real-world applications [^2] https://github.com/HRItdy/claw_machine.
Pipelines for Robot Systems Instruction Grounding Manipulation Generalization Adaptation for Deployment Visual inputs Multi-round dialogue CoT reasoning Morphology independent Skill sets Data Efficiency End-to-End VLA Models Wide range Data-hungry finetuning Modular VLM pipelines Controller- specific Cheap finetuning Multimodal LLMs Agents In-context learning
Table 1: Comparison of foundation model integration strategies in embodied robotic systems, highlighting differences in instruction grounding, manipulation generalization, and adaptation methods.
2 Foundation Model Integration for Language-Guided Robotics
Concerning how FMs are integrated into robot systems, we identified the following three types of integration strategies (Fig. 2). In the following, We briefly describe each strategy along with its respective advantages and limitations.
Figure 2: Foundation model integration strategies in language driven robots: (a) End-to-end VLA models, (b) Modular VLM pipelines, and (c) Multimodal LLM agents. Each strategy reflects a distinct interface between language, perception, and control.
2.1 End-to-End Vision-Language-Action
Definition.
Vision-Language-Action (VLA) models operate in an end-to-end manner, directly translating visual observations and natural language instructions into low-level actions without decoupled perception, language, and control modules (Fig. 2a). Two mainstream paradigms have emerged within this framework: auto-regressive and diffusion-based action generation. Through large-scale pretraining, these models acquire broad capabilities that support generalization across tasks. However, efficient adaptation to real-world settings remains a significant challenge.
Autoregressive VLA Models.
Autoregressive VLA models typically process language and visual inputs, employing various tokenization strategies to convert multimodal data into a unified latent space. Then the transformer-based decoder generates actions step-by-step in an autoregressive manner conditioned on the input context and previously generated actions, allowing structured action generation and planning.
Building on the previous groundbreaking models such as RT-1 Brohan et al. (2023), RT-2 Zitkovich et al. (2023) and VIMA Jiang et al. (2023), OpenVLA Kim et al. (2024) emerges as an important open-source method, combining a fine-tuned Llama 2 (7B) model with DinoV2 and SigLIP for visual tokenization, pretrained on the Open-X-Embodiment dataset Collaboration et al. (2024) consisting of 970k real-world robot demonstrations. TraceVLA Zheng et al. (2025) improves OpenVLA with visual trace prompting to enhance spatiotemporal awareness. Emma-X Sun et al. (2024) further refines dataset quality using a trajectory segmentation and Embodied Chain-of-Thought reasoning. This work is followed by NORA Hung et al. (2025), which uses Qwen-2.5-VL-3B as backbone and pretrained on the Open-X-Embodiment dataset. In parallel, other efforts integrate tactile sensing into robot perception Yang et al. (2024); Zhao et al. (2024), supporting more generic robot policies. Other recent developments include distilling spatial representations from VLMs, e.g., Gemini-Robitics Gemini Robotics Team (2025), and enriching training with additional features, such as 3D spatial relationships in SpatialVLA Qu et al. (2025), task-centric latent space in UniVLA Bu et al. (2025), and integrating multimodal understanding with action prediction in ChatVLA Zhou et al. (2025) and UP-VLA Zhang et al. (2025).
Diffusion-based VLA Models.
Diffusion-based VLA models formulate action generation as a denoising process over latent trajectories. Given a noisy version of a full action sequence, the model learns to recover the true trajectory conditioned on language and vision.
Diffusion Policy (DP) Chi et al. (2023) pioneered the use of diffusion model for visuomotor policy representation, laying the foundation for subsequent multimodal approaches. Octo Team et al. (2024) uses conditional diffusion decoding for action sequence prediction. Rather than lightweight diffusion heads, Reuss et al. (2024); Wen et al. (2024); Li et al. (2024b) use larger and more dedicated diffusion policy modules the as action decoder. DTP Fan et al. (2025) introduces a trajectory-level guidance to enhance diffusion-based planning. Recent works such as $\pi_{0}$ Black et al. (2024) and $\pi_{0.5}$ Intelligence et al. (2025) integrate a pretrained VLM with a flow-matching-based action expert to model action distributions. These models often utilize FAST Pertsch et al. (2025) for efficient compressed action tokenization. HybridVLA Liu et al. (2025) unifies diffusion and autoregressive action prediction within a single LLM-based framework. A growing focus is placed on adapting VLA models to diverse embodiments including bimanual manipulation and humanoid robots, such as RDT-1b Liu et al. (2024b), DexVLA Wen et al. (2025) and GR00T N1 Bjorck et al. (2025).
Strengths and Limitations.
VLA models provide a unified, end-to-end framework for robotic manipulation, (i) seamlessly integrating visual, language, and action modalities. Leveraging large-scale pretraining and fine-tuning, these models exhibit (ii) strong potential for generalizing across diverse manipulation tasks and robotic embodiments. However, their performance is constrained by the (iii) limited availability of high-quality, diverse robotic datasets. Pretraining can also introduce biases from training distributions, leading to (iv) degraded performance in out-of-distribution scenarios, such as novel tasks or different robotic embodiments. Therefore, despite their potential, further work is needed to enhance their robustness and generalization across real-world settings, for example, efficient adaptation using few-shot data.
2.2 Modular Vision–Language Pipelines
Definition.
In a modular vision-language pipeline (Fig. 2b), perception is handled by a specialist vision language model (VLM) that outputs symbolic scene information, typically grounded 2-D / 3-D bounding boxes, segmentation masks, or referring expression pointers. A downstream planner or policy module then consumes this structured representation to generate low-level actions. The language channel is therefore disentangled from motor control, allowing each module to be tuned independently, thus preserving the transparency and plug-and-play advantages of classical planning.
Representative systems.
Language-promptable specialist VLMs endow modular stacks with zero-shot semantics for various robotics pipelines. Bandyopadhyay et al. (2024) demonstrates an end-to-end sample collection robot system that uses GroundingDINO Liu et al. (2024a) to localize objects and refines each box with SAM Ravi et al. (2025) masks before passing them to classical grasp-and-place controllers, illustrating this paradigm’s practicality in real deployments. Werby et al. (2024) aggregates these modules into a floor–room–object hierarchy, showcasing their usage in long-horizon language-conditioned navigation across multi-story buildings.
Strength and Limitations.
Modular VLM pipelines strike a balance between transparency and adaptability, and delivers practical benefits: (i) interpretability—detections can be inspected; (ii) lightweight—the model parameters are usually around 100Msimilar-$\sim$600M, approximately 1% $\sim$ 6% the size of LLaMA 3.2 Vision 11B Grattafiori et al. (2024). On the other hand, it is limited at (i) interaction rigidness compared with more flexible multimodal LLMs, and (ii) pipeline brittleness where perception errors propagate without mitigation (Fig. 2b; Table 1). Their success hinges on robust open-vocabulary grounding—precisely the capability our Instruction Grounding case study stresses in Section 3.
2.3 Multimodal LLM Agents as Orchestrators
Definition.
Multimodal LLM agents place a large, tool-calling language model at the centre of the control loop (Fig. 2c). The LLM receives raw user utterances, selectively invokes vision tools (e.g., a detector or depth estimator) via function calls, reasons over their outputs in-context, and finally issues high-level action primitives to a low-level controller. The agent therefore acts as a cognitive hub that binds perception and control through natural language.
Representative Systems.
Multimodal LLMs are taking increasingly important roles in robotics. Gemini Robotics Gemini Robotics Team (2025) integrates perception, spatial reasoning, and trajectory synthesis into one Gemini-2.0 backbone Google DeepMind (2024), which serves as the embodied brain. Li et al. (2024c), in the same vibe as Gemini Robotics, leverages the inherent common sense and reasoning capabilities of these models by fine-tuning adapter modules through a chain-of-thought training paradigm. It endows the model with accurate pose prediction and precise manipulation abilities. These works collectively show the trend that the multimodal LLM shifts to the “cognitive hub” in robot systems. Glocker et al. (2025) build a modular agent-orchestration system for household object management robots. It utilize Llama 3.2 Vision Grattafiori et al. (2024) for open-vocabulary perception to facilitate creating grounded task plans, while the limitations of the multmodal LLM were not discussed.
Somewhat similar to our work, Li et al. (2024a) investigates the eligibility of Multimodal LLMs to serve as the “brain” for in-home robotics by providing a benchmark to compare models along the axes of perception, visual reasoning and task planing. Models like GPT-4V, Qwen-VL Bai et al. (2025) and DeepSeek-VL Lu et al. (2024) were included, but more recent releases were not covered—likely due to the fact that the field is moving fast with new models emerging in rapid succession.
Strengths and Limitations.
Multimodal LLM agents excel in (i) visual commonsense reasoning, leveraging extensive language priors to generalize to novel concepts beyond the reach of most specialist VLMs, and (ii) instruction following with support for fine-grained visual understanding and dynamic planning. Despite their expressive power, however, these models are (iii)  resource-intensive, posing challenges for deployment—particularly on mobile robotic platforms.
3 Case Studies on Instruction Grounding
Figure 3: Experimental setup for two case studies in a cluttered tabletop environment. The top row shows egocentric video data collected for the manipulation case study. The bottom row is an example setup for the instruction grounding task, including an annotated visual prompt paired with complex instructions in three forms: implicit, explicit with attributes and spatial references.
Natural language instruction grounding involves translating user intents into clear, actionable goals in a visual scene, which is a key capability for embodied AI Gemini Robotics Team (2025). Our case study offers empirical insights into the grounding performance of various models through the lens of challenging cross-modal disambiguation, and further examines the trade-offs introduced by model sizes and quantization—providing practical suggestions for efficient deployment.
Benchmark Dataset.
To minimize the impact of vision priors on measured performance, we design benchmarking scenarios using household objects placed on a tabletop. These objects are commonly represented in the training datasets of the foundation models, and the tabletop setup features minimal variation in lighting and camera angles—ensuring that the evaluation primarily reflects grounding capabilities.
We curated a new Instruction Grounding benchmark dataset. In images containing multiple household objects, each object is tagged with a number as the visual prompt, and each image is paired with language instructions crafted to test visual commonsense and cross-modal disambiguation concerning attribute or spatial relationships – for “pick up the red-capped marker,” the color must be used to select one among a few markers; whereas “grasp the cup in front of the screwdriver” requires reasoning over spatial relations (Fig. 3; Appendix).
Language ambiguity often leads to execution failure in an embodied system. By comparing specialist VLMs and multimodal LLMs, we reveal their concrete failure modes that further inform our design implications in Sec. 5.
Model Easy Medium Hard Avg Specialist VLMs GroundingDINO-86M 0.518 0.357 0.349 0.408 GroundingDINO-145M 0.443 0.320 0.355 0.372 Closed-source Multimodal LLMs Gemini2.5-Pro-Exp 0.904 0.765 0.793 0.821 Gemini2.0-Flash 0.884 0.738 0.678 0.767 GPT-4.5 0.837 0.723 0.739 0.766 GPT-4o 0.814 0.745 0.683 0.747 GPT-4o-mini 0.803 0.722 0.604 0.710 o4-mini 0.721 0.769 0.710 0.733 GPT-4V 0.470 0.476 0.467 0.471 Open-source Multimodal LLMs Llama-3.2V-90B 0.722 0.701 0.657 0.693 Llama-3.2V-11B 0.583 0.569 0.547 0.566 Llama-4-Maverick 0.698 0.576 0.634 0.636 Llama-4-Scout 0.776 0.615 0.624 0.672 Qwen2-VL-72B 0.686 0.614 0.558 0.619 Gemma-3-27B 0.452 0.384 0.267 0.368 DS-Janus-Pro-7B 0.444 0.330 0.317 0.364 Phi-3.5-Vision-4.2B 0.291 0.357 0.205 0.284
Table 2: Object grounding performance of specialist VLMs and multimodal LLMs (closed-source and open-source) across varying scene complexity levels. Models are evaluated on easy, medium, and hard cluttered scenes, with macro accuracy reported.
Figure 4: Performance of complex instruction grounding across VLM–LLM pipelines and end-to-end multimodal LLMs. Macro accuracy is reported across instruction types—implicit, attribute-based, and relationship-based. Subfigures show (a) proprietary models and (b) open-source models along with their Int4-quantized variants.
Zero-Shot Object Grounding.
We begin with a foundational question for instruction grounding: Can FMs accurately recognize objects in cluttered open scenes? Table 2 presents the performance hierarchy for specialist VLMs and a range of multimodal LLMs, serving as a basis for deeper analysis of ambiguity resolution in later sections.
  • Despite their popularity in modular pipelines, GroundingDINO achieve only 35–41% accuracy, as it struggles with featureless objects, e.g. ‘the can’. Moreover, it is brittle in open scenes, e.g. a ‘screwdriver’ is constantly recognized as a ‘marker’, which instead is an easy case for multimodal LLMs which embodied large volume of visual commonsense.
  • Gemini 2.5-Pro takes the first place with 0.82, followed by Gemini 2.0-Flash and GPT-4.5. Open-source systems still trail the proprietary tier. Llama 3.2-Vision 90B reaches about 84% of Gemini 2.5’s score, while the more recent Llama 4 releases did not outperform it.
  • Smaller community models (Gemma-27B, Phi-Vision) fall below the specialist-threshold, suggesting that they are still inadequate for fine-grained grounding in cluttered scenarios.
  • Last, when compute is a bottleneck, GPT-4o-mini (0.71) and Llama 3.2-Vision 11B (0.57) provide the best speed–accuracy trade-off, delivering decent performance without incurring heavy memory footprint or high API costs.
Zero-Shot Complex Instruction Grounding.
This task is framed as a multiple-choice problem, where the model is asked to select the correct object index in a cluttered scene based on three types of natural language instructions: implicit, attribute-based, and relationship-based—each type probes a distinct grounding challenge. We evaluate a series of multimodal LLMs, using a modular VLM–LLM pipeline as a baseline. In this pipeline, the LLM parses the instruction to infer likely targets, queries GroundingDINO to detect candidate objects, and selects from the detected boxes—essentially guessing without directly perceiving the scene.
  • Implicit Instruction Grounding. Instructions like “I need a tool to tighten the screws” only refer to the target object implicitly, and the model needs to infer the target object using its common sense priors. For such instructions, the modular VLM–LLM pipeline struggles to select a screwdriver, lacking embedded affordance reasoning. In contrast, multimodal LLMs perform well, reflecting strong visual commonsense. GPT-4.5 demonstrates exceptional performance (0.94), though its high inference cost—20×20$\times$20 × that of Gemini 2.5 makes it cost-prohibitive for most applications.
  • Relational Reasoning Remains Challenging. This category requires resolving referential ambiguity through implicit chain-of-thought reasoning: grounding objects, modeling spatial relationships, and disambiguating targets (e.g., identifying the correct mug among many based on “next to something”). Accuracy drops significantly nearly across all models. Only Gemini 2.5-Pro and o4-mini achieve accuracy above 0.80—the former likely benefits from embodied training data, while the latter demonstrates strong reasoning capabilities. Notably, o4-mini is a medium-sized model, yet it outperforms larger models like GPT-4.5 on relational instructions—suggesting that structured reasoning may help close, or even overcome the performance gap brought by different model scales.
  • Instruction-Dependent Quantization Effects. INT4 quantization reduces the model size by over 70%, making it an attractive choice for deployment. In Llama 3.2 Vision, we observe that it disproportionately impacts implicit and relational instruction grounding, indicated by the relative accuracy drop of $14\%-17\%$, while attribute grounding is more robust with only 4% loss. Despite reduced precision, quantized 11B models offer a speed–accuracy balance for low-resource settings. Our findings underscore the need for fine-grained quantization strategies that preserve the most important high-level reasoning capabilities under resource constraints.
Models LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
OpenVLA finetuned 84.7 88.4 79.2 53.7 76.5
$\boldsymbol{\pi_{0}}$finetuned 96.8 98.8 95.8 85.2 94.15
[…]
SpatialVLA finetuned-AC 88.2 89.9 78.6 55.5 78.1
NORA-finetuned 85.6 87.8 77.0 45.0 73.9
NORA-finetuned-AC 85.6 89.4 80.0 63.0 79.5
NORA-Long-finetuned 92.2 95.4 89.4 74.6 87.9
Table 3: Success rates (%) on the LIBERO Simulation Benchmark across four task suites, each evaluated over 500 trials. Results for SpatialVLA are from Qu et al. (2025); Results for $\pi_{0}$ are from Black et al. (2024), using pretrained models on LIBERO benchmarks. “AC” denotes the use of action chunking. The comparison in the Appendix highlights its impact on performance. The finetuned $\pi_{0}$ model achieves the highest performance.
4 Case Studies on Robotic Manipulation
Now we shift the focus to skill adaptation. In an ideal deployment scenario, a pretrained VLA—already endowed with broad visuomotor skills—should be retargeted to a new manipulation task with minimal data and fast convergence. We use fine-tuning, the standard practice for adaptation, as a probing lever to evaluate how the state-of-the-art VLA models adapt to new tasks and deployment conditions.
Given the scale of VLAs, we compare partial fine-tuning, which leverages our benchmark dataset (Appendix) and its inherent distribution bias to study convergence behavior, and full fine-tuning, which uses large-scale datasets to minimize the training loss. Our evaluation focuses on three key aspects: (i) training dynamics—how quickly and smoothly training converges; (ii) generalization—how well the resulting policies perform on various tasks; and (iii) robustness—how well the resulting policies handle environmental distractors. Our experiments highlight the performance of VLA models in different settings, offering practical suggestions for practitioners who have to adapt large VLAs under tight data, time and compute budgets.
Skill Adaptation Performance.
Our fine-tuning process consists of two stages: (1) To assess convergence behavior under distribution shift, we collected a custom dataset (see Appendix for details) with a distribution bias relative to common pretraining datasets included in the Open-X-embodiment Collaboration et al. (2024) and LIBERO datasets Liu et al. (2023). We used it to partially fine-tune several recent VLA models and trained Diffusion Policy (DP) and Action Chunking Transformer (ACT) from scratch. The results are shown in Fig. 5; (2) For full fine-tuning, we leveraged larger benchmark datasets, Open-X-embodiment and LIBERO, to fully fine-tune RT-1, OpenVLA, SpatialVLA and NORA, and compared their performance. The results are shown in Fig. 6.
  • Partial Fine-tuning. Through the experiments we observe that DP and ACT exhibit high stability with low variance during training. In contrast, generalist models such as OpenVLA and $\pi_{0}$ require significantly more training iterations to attain comparable accuracy and exhibit greater variance, which can be attributed to their large model capacity. Notably, although DP achieves lower loss by fitting directly to noise, it still demands more training steps to generate coherent actions, even after loss convergence.
  • Full Fine-tuning. These fully fine-tuned VLA models are evaluated on three tasks: (1) out-of-distribution (OOD) object manipulation, (2)  spatial relationship reasoning and (3) multi-object pick-and-place tasks. In task (1), both NORA and OpenVLA succeed, while SpatialVLA fails due to incorrect affordance point estimation. In task (2), NORA correctly follows instructions, while OpenVLA fails and SpatialVLA exhibits unstable performance. In task (3), NORA achieves successful execution while other models fail to complete the task reliably.
Sim2Real Adaption Performance.
We compare model performance on simulation benchmarks and real robot deployments. A significant drop in performance is observed during transfer from simulation to the real world (Table 3; Appendix Table 6). The simulation benchmark includes 30 procedurally-generated, disentangled tasks requiring nuanced spatial reasoning (LIBERO-Spatial), object understanding (LIBERO-Object), and goal interpretation (LIBERO-Goal), as well as 10 long-horizon, entangled tasks (LIBERO-Long).
Robustness to Perturbations.
To evaluate robustness, we introduced distractor objects into the environment. As shown in Table 8, both OpenVLA and NORA exhibit substantial performance degradation in the presence of these perturbations, highlighting their sensitivity to novel conditions.
Figure 5: Results for partial fine-tuning of VLA models including OpenVLA and $\pi_{0}$, alongside results from training Diffusion Policy and ACT (Action Chunking Transformer) from scratch on our dataset. VLA models require more training epochs to converge and exhibit higher variance in performance.
Key Takeaways.
Current VLA models still face significant limitations in the following areas:
  • Adaptation and Generalization. A generic robotic policy is expected to quickly adapt to datasets with distributional shifts. However, according to the partial fine-tuning results, due to the large model capacities and the limited size of task-specific datasets, these VLA models failed to achieve fast adaptation. While full fine-tuning offers improved performance, it requires extensive data and long training time, which are impractical for many real-world scenarios.
  • Robustness. Robustness to distribution shifts (without finetuning) is a critical challenge. Results reveal substantial performance degradation both when encountering unseen objects and during sim-to-real transfer, highlighting the fragility of current VLA models in dynamic and unpredictable environments.
These findings suggest that while VLA models hold promise, they have limitations in data efficiency, adaptation speed, and robustness to make them reliable for real-world robotic applications.
Figure 6: Success rates of fully fine-tuned VLA models on out-of-distribution object manipulation (OOD Object), spatial relationship reasoning (Spatial) and multi-object pick-and-place (Multiple) tasks. NORA achieves the highest performance.
5 Constraints and Future Directions
Despite the promise of foundation models for enabling embodied agents to perform daily tasks, the following critical constraints still hinder their widespread deployment:
Data Scarcity.
In contrast to natural language datasets, which are readily sourced from internet, robotic datasets are significantly more expensive due to high hardware costs and intensive labor during data acquisition. A promising direction for future research is developing more data-efficient models. In addition, exploring high-fidelity simulation environments and developing robust sim-to-real transfer techniques could mitigate data scarcity.
Limited Generalization Capability.
A key limitation of current VLA models is their limited ability to generalize to out-of-distribution concepts that were not well-represented during training, which is a consequence of the aforementioned data scarcity. Many models depend heavily on large-scale paired datasets, which often exhibit biases and limited diversity in aspects such as camera viewpoints, lighting conditions and specific robotic embodiments. This results in fragile performance when deployed in real-world or domain-specific scenarios. Furthermore, these models struggle with fine-grained spatial reasoning and temporal understanding, hindering them from accurately aligning language with complex visual scenes or dynamic events.
Efficient Inference.
Deploying large-scale models on robotic platforms introduces significant computational challenges, primarily in inference speed and GPU RAMs. This underscores the importance of smaller models that can efficiently generate actions without significant performance degradation. This issue is particularly pronounced in autoregressive models, and diffusion models are less affected.
Limitations
When evaluating the generalization capabilities of Vision-Language Alignment (VLA) models, this paper primarily focuses on different tasks without extensively addressing their generalization across varying robot morphologies. Although relatively few works have specifically targeted this aspect, it remains a significant challenge in deploying VLA models in real-world applications. Robots with different morphologies, such as bimanual manipulators, humanoid robots, and autonomous vehicles, require distinct operational protocols and safety considerations. The absence of a generic robot policy that can adapt seamlessly across diverse morphologies limits the practical generalization potential of VLA models and hinders their deployment as universal robotic policies.
In addition, this paper does not explore the ability of VLA models to ground instructions involving open-ended or ambiguous commands. Current VLA models are largely trained on curated datasets, which allow them to learn mappings from instructions to specific actions. However, this reliance constrains their ability to truly understand instructions at the semantic level. As a result, when facing out-of-distribution or vague commands, these models often struggle to infer reasonable actions. Addressing this limitation will require integrating more advanced instruction-understanding modules into the VLA pipeline to improve their robustness in handling ambiguous or under-specified input.
Appendix
Appendix A Benchmark Dataset
A.1 Cluttered Tabletop Manipulation Dataset
To evaluate the finetuning behavior of various VLA models under distribution shift, we constructed a custom cluttered tabletop environment using a UR5 robotic arm with a wrist-mounted RealSense RGB-D camera. This setup differs from all existing configurations in the Open-X-Embodiment dataset. Demonstrations for a screwdriver-picking task—amid distractor objects—were collected via teleoperation using a SpaceMouse device. In total, we gathered 163 demonstration episodes [^3] https://huggingface.co/datasets/bittdy/pick_screw. Each episode began with a randomized initial robot pose, followed by an attempt to grasp the target screwdriver.
A.2 Complex Instruction Grounding Dataset
We curated an evaluation dataset for the complex instruction grounding task in cluttered scenes [^4] https://github.com/xiuchao/InstructionGrounding. Thirty images were sampled from the action sequences and subsequently categorized based on the number of objects: Easy (<15), Medium (15$\sim$20), and Hard (>20). Objects in the visual scenes were manually annotated using visual prompts and paired with various instructions. The spatial relationship words were illustrate in Table 4.
Words Positional:                 left, right, between, beside, near, far, front, behind Directional:           aligned with, perpendicular to Instruction hand over the screwdriver [on the left of] the red ball. pass me the screwdriver [aligned with] the marker.
Table 4: Template words and corresponding examples of generated relation-based instructions for case studies.
For the complex instruction grounding task, including an annotated visual prompt paired with complex instructions in three forms: implicit, explicit with attributes and spatial references. Additionally, the dataset includes multi-turn questions that refer to more than one object, enabling foundation models to ask clarifying questions to identify the correct object.
  • Implicit Instructions. Here, objects are not explicitly mentioned by name or attributes but are instead described by their functions. This category evaluates the VLMs’ ability to infer the correct object based on its use. For example, the dataset includes instructions referring to objects like scissors, screwdrivers, and rulers based on their respective functions.
    Figure 7: Example of visual prompts
  • Explicit Attributes. In this category, instructions prompt VLMs to identify objects belonging to a category with multiple instances, where each instance can be uniquely identified by explicitly mentioned attributes. In Fig. 7, the beige mug and the gray mug are included because they are unique when described with attributes. However, objects like the black mug or scissors are excluded. This is because there are two identical black mugs, making them non-unique, and there is only one pair of scissors, which does not require attributes for identification.
  • Explicit Relationships. In this category, instructions describe objects by their spatial relationships to other objects in the image. We ensure that each referenced object is unique within the image. For example, the measuring cup to the right of the screwdriver uniquely identifies the object. These instructions are designed to test the VLMs’ ability to comprehend and resolve location-based relationships.
  • Multi-Referent Instructions. This category includes instructions that correspond to multiple valid objects in the scene. For example, in Fig. 7, an instruction like “give me a mug” may refer to several similar items. In such cases, we annotate the data with all candidate object indices, e.g., [2, 7, 18], indicating the set of plausible referents.
Figure 8: Dataset breakdown by Instruction Types.
Model Easy Medium Hard
im attr rel im attr rel im attr rel
VLM-LLM 0.050 0.516 0.131 0.010 0.336 0.186 0.000 0.318 0.174
Gemini-2.5-Pro 0.778 0.889 0.830 0.847 0.814 0.815 0.985 0.784 0.858
Gemini-2.0 0.833 0.889 0.774 0.819 0.721 0.642 1.000 0.668 0.469
GPT-4.5 0.944 0.889 0.894 0.917 0.838 0.722 0.958 0.719 0.698
GPT-4o 0.850 1.000 0.778 0.819 0.948 0.680 0.901 0.697 0.469
o4 0.950 1.000 0.907 0.847 0.988 0.837 0.917 0.809 0.804
4o-mini 0.750 0.717 0.550 0.764 0.771 0.596 0.750 0.382 0.248
GPT-4V 0.650 0.750 0.598 0.750 0.737 0.662 0.625 0.417 0.455
Qwen2-VL 0.800 0.917 0.830 0.792 0.756 0.738 0.875 0.700 0.529
LLaMA 3.2 Vision 90B 0.750 0.850 0.704 0.708 0.853 0.711 0.875 0.491 0.521
LLaMA 3.2 Vision 90B-Q4 0.800 0.667 0.598 0.625 0.719 0.554 0.542 0.464 0.300
LLaMA 3.2 Vision 11B 0.650 0.667 0.631 0.764 0.710 0.556 0.833 0.536 0.342
LLaMA 3.2 Vision 11B-Q4 0.650 0.567 0.502 0.694 0.757 0.555 0.542 0.498 0.450
Table 5: Performance on the complex instruction grounding task. Abbreviations: im denotes implicit instruction, attr denotes attribute-based instruction, and rel denotes relation-based instruction.
A human-in-the-loop process was employed to ensure high-quality data collection.
  • Initial Object Identification: We used GPT-4o to identify objects in an image and referring them by type, explicit attributes, and detailed location relations.
  • Human Verification. The authors of this paper reviewed and modified the outputs to ensure their correctness.
  • Instruction Generation. After verification, GPT-4 was tasked with generating simple, clear instructions for different objects.
  • Final Review. These instructions underwent another round of verification to ensure clarity and accuracy.
This high-quality dataset consisting of 473 instructions, with a detailed breakdown of each instruction type presented in Fig. 8.
Figure 9: Examples of Instruction Grounding. (a) “the marker on the left”, (b) “the marker aligned with the ruler”.
Figure 10: Examples of Object Grounding. (a) “ball”, (b) “screwdriver”, (c) “marker pens”, (d) “blue ball”.
Appendix B Grounding Experiments
B.1 Complex Instruction Grounding for Goal Specification
Cross-modal Disambiguation represents a particularly challenging component of goal specification. To quantify the model capability in this dimension, we employed attribute-based and relative relationship instructions to uniquely identify a target among multiple candidates. The goal specification task is formulated as follows:
Given a visual input $I\in\mathbb{R}^{H\times W\times 3}$ and an instruction $t$, the objective is to predict the target object according to $o^{*}=\arg\max_{o\in\mathcal{O}}P(o\,|\,I,t;\theta_{FM})$, where $o^{\star}\in\mathcal{O}$ and $\mathcal{O}$ denotes the set of candidate objects. Models are evaluated using macro-average accuracy metric.
B.2 Failure Cases of Specialist VLM Pipeline
Grounding DINO, despite popular for zero-shot detection, is not robust in open scenes. It successfully detected “blue ball” while failed to detect “ball”, indicating its reliance on visual features. Similarly, featureless metal cans pose a great challenge for Grounding DINO, which were almost omitted in the detection results.
For complex instruction grounding, Grounding DINO and GPT-4 were chained together to “guess” the target by the LLM based on the candidate bounding boxes. The failure cases were illustrated in the Fig. 9 and Fig. 10.
B.3 Multimodal LLMs Performance
The performance of Multimodal LLMs on complex grounding across Easy, Medium and Hard groups are shown in Table 5.
Appendix C Manipulation Experiments
OpenVLA and $\pi_{0}$ were partially fine-tuned on this dataset, while Diffusion Policy (DP) and Action Chunking Transformer (ACT) were trained from scratch. Due to the limited size of our custom dataset, full fine-tuning of RT-1, OpenVLA, SpatialVLA, and NORA was performed using the Open-X-embodiment and LIBERO datasets. We evaluated model performance on both a real-world WidowX robotic platform and the LIBERO simulation benchmark.
C.1 Fine-tuning details for VLA
Partial fine-tuning was conducted on a single NVIDIA A6000 GPU (48 GB VRAM) over a period of three days. To ensure a fair comparison, a batch size of 1 was used across all models. The results are presented in Fig. 5.
Full fine-tuning of RT-1, OpenVLA, SpatialVLA, and NORA was conducted on a compute node equipped with 8$\times$×H100 GPUs. The fine-tuned models were evaluated on 9 diverse real-world manipulation tasks, as shown in Fig. 11. Success rates are summarized in Table 6, demonstrating NORA’s superior policy generation capabilities across three task categories: out-of-distribution object grasping, spatial reasoning, and multi-object manipulation.
Figure 11: Real-world robot environments and task setups. We evaluate these models across 9 diverse tasks to assess its instruction understanding, spatial reasoning, and multi-task motion planning capabilities.
Category Task RT-1 OpenVLA SpatialVLA NORA
Multiple objects Put the red bottle and the hamburger in the pan 0 20 0 40
Put the carrot and hotdog in pot 0 0 0 30
Put the corn and carrot in the pan 0 30 0 30
OOD object put carrot in pot 0 80 20 90
Put banana in pot 1 40 0 90
Put the blue cube on the plate 0 50 0 70
Spatial Put the pink toy at the right corner 0 60 30 60
Put the blue cube on the right plate 0 30 0 20
Move the banana close to the pan 30 50 50 80
Average 4.4 40 11.1 56.7
Table 6: Task performance comparison across different categories and models.
C.2 Impact of Action Chunking
C.2.1 Action Chunking Performs on WidowX.
To investigate the effectiveness of action chunking, we selected NORA-LONG and SpatialVLA for evaluation. Tasks were chosen from three categories: (1) “put the carrot in the pot,” (2) “put the red bottle and hamburger in the pot,” and (3) “put the pink toy at the right corner.” In initial experiments, all predicted actions (5 actions for NORA-LONG, 4 actions for SpatialVLA) were executed sequentially without replanning. This frequently caused the WidowX robot to crash into the environment due to the accumulation of overly large movements.
Subsequently, we modified the execution policy to only perform the first action in each predicted chunk. This adjustment resolved the collision issue and NORA -LONG achieved an 80% success rate on the “put the carrot in the pot” task. However, on multi-object pick-and-place tasks, NORA-LONG consistently stopped after placing the first object, resulting in a 0% final success rate. For the spatial reasoning task, NORA-LONG achieved a 70% success rate on “put the pink toy at the right corner.”
C.2.2 Action chunking improves performance in simulation.
We hypothesize that action chunking is more effective at higher control frequencies. For example, Diffusion Policy generates commands at 10 Hz, which are then interpolated to 125 Hz for execution. Similarly, OpenVLA-OFT+ employs action chunking and shows improved performance in real-world ALOHA tasks, which run at 25 Hz.
Since our real robotic platforms do not support high-frequency control, we tested this hypothesis in the LIBERO simulation environment (20 Hz). We fine-tuned both NORA and NORA-LONG on this benchmark with an action chunk size of 5, producing two variants: NORA-finetuned-AC and NORA-Long-finetuned.
Results show that NORA-finetuned-AC significantly outperforms NORA-finetuned across all LIBERO benchmarks, with a higher average success rate. Notably, NORA-Long-finetuned outperforms all baseline models (see Table 3), highlighting the benefits of pretraining with action chunking and its transferability to long-horizon tasks. However, it is important to note that LIBERO is a simulation environment and may not reflect real-world performance at high control frequencies.
Figure 12: Comparison of tasks with and without distraction.
TASK OpenVLA NORA
without Distraction
put carrot in pot 8 9
put banana in pot 4 9
put the blue cube on the plate 5 7
with Distraction
put carrot in pot 6 8
put banana in pot 6 4
put the blue cube on the plate 3 5
Table 7: Comparison of task performance between OpenVLA and NORA under conditions with and without distraction. Each value denotes the number of successful executions out of 10 trials.
C.3 Robustness to Disturbance
To evaluate robustness, we selected three straightforward tasks (shown in Fig. 12) and introduced distractor objects into the environment. Initially, both OpenVLA and NORA performed well. However, their success rates declined significantly with the introduction of distractions. This highlights the sensitivity of current VLA models to out-of-distribution disturbances. The average success rates across the three tasks are presented in Table 8, while the detailed number of successful executions out of 10 trials is summarized in Table 7.
Table 8: Average Success Rate (%) without (w/o) and with (w/) Distractors
Model w/o Distractors w/ Distractors
OpenVLA 56.7 50
NORA 83.3 56.7
Appendix D Modular Claw Machine Prototype
To facilitate the evaluation of different VLMs in robotic manipulation, we developed a voice-controlled testbed using a UR5 robotic arm [^5] https://github.com/HRItdy/claw_machine. The system architecture, shown in Fig. 13, comprises the following five modules:
  • Speech Transcription: Powered by Microsoft Azure’s speech recognition service.
  • Task Decomposition: Based on GPT-3.5 and GPT-4 using prompting paradigms adapted from ChatGPT for Robotics.
  • Object Detection: Utilizes GroundingDINO and OWL-ViT for object detection.
  • Object Segmentation: Employs Segment Anything Model (SAM) and FastSAM for segmenting detected objects.
  • Manipulation: Low-level actions are generated by GraspAnything or GraspNet.
This modular testbed enables rapid integration and benchmarking of different models within a real robotic system.
Figure 13: The system architecture of the testbed for VLMs.