### v2

# From Grounding to Manipulation: Case Studies of Foundation Model Integration in Embodied Robotic Systems

Abstract
Foundation models (FMs) are increasingly applied to bridge language and action in embodied agents, yet the operational characteristics of different integration strategies remain under-explored—especially for complex instruction following and versatile action generation in changing environments. We investigate three paradigms for robotic systems: end-to-end vision-language-action models (VLAs) that implicitly unify perception and planning, and modular pipelines using either vision-language models (VLMs) or multimodal large language models (MLLMs). Two case studies frame the comparison: instruction grounding, which probs fine-grained language understanding and cross-modal disambiguation; and object manipulation, which targets skill transfer via VLA finetuning. Our experiments reveal trade-offs in system scale, generalization and data efficiency. These findings indicate design lessons for language-driven physical agents and point to challenges and opportunities for FM-powered robotics in real-world conditions.
From Grounding to Manipulation: Case Studies of Foundation Model Integration in Embodied Robotic Systems
Xiuchao Sui1thanks: Equal contribution. Corresponding authors. {Sui_Xiuchao, Tian_Daiying}@a-star.edu.sg  and Daiying Tian [^1] footnotemark: 1  and Qi Sun 2  and Ruirui Chen1  and Dongkyu Choi 1and Kenneth Kwok 1and Soujanya Poria 2 1IHPC, Agency for Science, Technology and Research, Singapore 2Nanyang Technological University, Singapore
1 Introduction
Natural language is emerging as a universal interface for embodied robotics. Advances in foundation models (FMs) enable robots to follow free-form instructions across perception, reasoning, and motor control, offering the promise of language-grounded autonomy. These models include vision–language models (VLMs) Liu et al. (2024a); Ravi et al. (2025); Ren et al. (2024); Li et al. (2023), multimodal large language models (MLLMs) Grattafiori et al. (2024); Bai et al. (2025); Lu et al. (2024), and vision–language–action (VLA) models Kim et al. (2024); Zheng et al. (2025); Qu et al. (2025); Bu et al. (2025).
However, realizing the promise of language-grounded autonomy in deployable systems remains highly challenging. Robots must (i) map ambiguous instructions to the physical world (instruction grounding), (ii) execute reliably across novel objects, scenes, and morphologies (generalizable execution), and (iii) achieve these goals with limited data (efficient adaptation). How effectively different FM integration strategies address these competing requirements remains under-explored (Fig. 1).
Figure 1: Key challenges of FMs in embodied robotics.
This work offers an empirical study of three integration paradigms: end-to-end VLAs mapping language and vision to actions, MLLM agents orchestrating perception and control, and modular VLM pipelines pairing perception-specialist FMs with task-specific planners (Fig.2; Table1).
We evaluate these paradigms through tabletop case studies that expose their complementary strengths and limitations. Two task categories are considered: complex instruction grounding, probing fine-grained understanding and cross-modal disambiguation (Sec.3); and object manipulation, assessing skill transfer after VLA fine-tuning under distribution shifts, complemented by comparative and ablation analyses (Sec.4).
Our instruction grounding experiments reveal distinct trade-offs across integration strategies. VLM pipelines emphasize interpretability and efficiency, but fall short of peak performance. They struggle with complex instruction grounding yet achieve moderate object grounding—using less than 1% of the parameters required by MLLMs. In contrast, MLLMs generalize better on complex instructions but incur substantially higher inference costs, with smaller reasoning-focused models at times outperforming larger ones. VLAs, with tightly coupled perception-to-action pathways, streamline action generation but make it difficult to isolate and assess grounding performance. To probe this gap, we evaluate their perception heads as reference points, which also surfaces open questions for VLA architecture design.
While MLLMs perform well on grounding tasks, their scale limits practical deployment on robotic platforms. Quantization is often adopted as a simple remedy, yet instruction-dependent behaviors emerge, with certain reasoning capabilities degrading disproportionately under naïve compression. This underscores the need for fine-grained quantization strategies. Taken together, our findings clarify the trade-offs between model scale and performance and provide practical guidance for building robotic systems under real-world constraints.
Our object manipulation experiments investigate the skill-adaptation capabilities of VLAs in both real-world and simulated settings. We examine how these models transfer manipulation skills under distribution shifts, evaluating their training dynamics, generalization, and robustness. Results reveal fragile training and slow adaptation in generalist policies, such as OpenVLA Kim et al. (2024), compared to task-specific counterparts, underscoring the challenges of real-world deployment.”
In summary, our main contributions are:
  • We analyse three FM integration paradigms on shared embodied tasks designed to probe the capabilities and trade-offs of FMs.
  • We release a dataset and code [^1] https://github.com/xiuchao/InstructionGrounding for evaluating instruction grounding and object manipulation, covering cross-modal reasoning and skill adaptation under varied layouts.
  • We providing timely insights into state-of-the-art VLAs and MLLMs, investigating their capabilities and failure modes and distilling practical trade-offs to guide practitioners in selecting FM stacks for language-driven embodied agents.
  • We also release a complete end-to-end claw-machine robot system as a real-world FM integration demo [^2] https://github.com/HRItdy/claw_machine.
Pipelines for Robot Systems Instruction Grounding Manipulation Generalization Adaptation for Deployment Visual inputs Multi-round dialogue CoT reasoning Morphology independent Skill sets Data Efficiency End-to-End VLA models Wide range Data-hungry finetuning Modular VLM pipelines Controller- specific Cheap finetuning Multimodal LLMs agents In-context learning
Table 1: Comparison of foundation model integration strategies in embodied robotic systems, highlighting differences in instruction grounding, manipulation generalization, and adaptation methods.
2 Foundation Model Integration for Language-Guided Robotics
Concerning how FMs are integrated into robot systems, we identified the following three types of integration strategies (Fig. 2). In the following, We briefly describe each strategy along with its respective advantages and limitations.
Figure 2: Three FM integration strategies for embodied robotics, highlighting distinct interfaces between language, perception, and control.
2.1 End-to-End VLA Models
Definition.
VLAs operate in an end-to-end manner, directly translating visual observations and natural language instructions into low-level actions without decoupled perception, language, and control modules (Fig. 2a). Two mainstream paradigms have emerged within this framework: auto-regressive and diffusion-based action generation. Through large-scale pretraining, these models acquire broad capabilities that support generalization across tasks. However, efficient adaptation to real-world settings remains a significant challenge.
Autoregressive Models.
These models generate actions step by step, with each action conditioned on the current perceptual input and previously generated outputs. These models typically encode visual observations and language into a shared latent space, then employ a transformer-based decoder to autoregressively predict low-level control tokens, such as joint angles, end-effector displacements and gripper states, which are typically obtained by discretizing continuous motor signals into token sequences.
RT-1 Brohan et al. (2023) and RT-2 Zitkovich et al. (2023) demonstrated the effectiveness of large-scale pretraining for task generalization. Building on these scaling successes, OpenVLA Kim et al. (2024) became a landmark open-source effort, combining Llama-2 with DinoV2 and SigLIP, trained on the large-scale Open-X-Embodiment dataset Collaboration et al. (2024). Subsequent works such as Emma-X Sun et al. (2024), NORA Hung et al. (2025) and TraceVLA Zheng et al. (2025), further advanced this line through dataset quality, stronger backbones and improved spatiotemporal prompting.
Subsequent models further incorporate richer modalities such as tactile sensing Yang et al. (2024); Zhao et al. (2024), structural priors including spatial representations distilled from VLMs Gemini Robotics Team (2025) and 3D spatial relationships in SpatialVLA Qu et al. (2025), and refinements to the embodied latent space, as in UniVLA Bu et al. (2025), which learns task-specific latent representations.
Diffusion-based Models.
They generate actions by progressively denoising trajectories from noise, modeling the distribution of future actions as a whole rather than step by step. Like autoregressive VLAs, these models encode visual observations and language into a latent representation, but instead apply a diffusion process that iteratively refines action sequences into feasible trajectories. Diffusion-based models trade inference speed for greater trajectory coherence.
Diffusion Policy Chi et al. (2023) pioneered the use of diffusion models for visuomotor policy learning, showing improved global consistency and robustness compared to autoregressive baselines. Octo Team et al. (2024) applies conditional diffusion decoding for action sequence prediction. Subsequent models further scales diffusion heads into larger, dedicated policy modules Reuss et al. (2024); Wen et al. (2024); Li et al. (2024b), and integrate trajectory-level guidance to improve long-horizon planning Fan et al. (2025). Notable models like $\pi_{0}$ Black et al. (2024) combine pretrained MLLM (PaliGemma-3B) with flow-matching-based action experts to produce continuous, high-frequency control across robot embodiments. $\pi_{0.5}$ Intelligence et al. (2025) builds on this by co-training across diverse environments and data modalities to improve generalization to new settings.
Beyond architectural advances, diffusion-based models increasingly targets diverse embodiments, extending from single-arm manipulation to bimanual systems and humanoid robots. Representative examples include RDT-1B Liu et al. (2024b), DexVLA Wen et al. (2025), and GR00T N1 Bjorck et al. (2025), which demonstrate the scalability of FM-driven action generation to more complex morphologies.
Strengths and Limitations.
Leveraging large-scale pretraining, VLAs hold the potential to generalize across diverse manipulation tasks and robot morphologies. Yet progress is constrained by the scarcity of high-quality, diverse robotic datasets, which limits both scale and coverage. Pretraining can also propagate biases from training distributions, causing degraded performance on novel tasks, in unseen environments, or with unfamiliar embodiments. Despite their promise, VLAs remain brittle in real-world deployment, as highlighted in our Object Manipulation case study (Section 4), underscoring the need for better data curation, bias mitigation, and efficient adaptation strategies such as few-shot and continual learning.
2.2 Modular VLM Pipelines
Definition.
In this paradigm (Fig. 2b), perception is handled by a specialist VLM that outputs symbolic scene information, typically grounded 2D/3D bounding boxes, segmentation masks, or referring expression pointers. A downstream planner or policy module then consumes this structured representation to generate low-level actions. The language channel is therefore disentangled from motor control, allowing each module to be tuned independently, thus preserving the transparency and plug-and-play advantages of classical planning.
Representative systems.
Language-promptable specialist VLMs endow modular stacks with zero-shot semantics for various robotics pipelines. Bandyopadhyay et al. (2024) demonstrates an end-to-end sample collection robot system that uses GroundingDINO Liu et al. (2024a) to localize objects and refines each box with SAM Ravi et al. (2025) masks before passing them to classical grasp-and-place controllers, illustrating this paradigm’s practicality in real deployments. Werby et al. (2024) aggregates these modules into a floor–room–object hierarchy, showcasing their usage in long-horizon language-conditioned navigation across multi-story buildings.
Strength and Limitations.
[…]
2.3 Multimodal LLM Agents as Orchestrators
Definition.
In this paradigm (Fig. 2c), MLLMs take raw user utterances, selectively invoke vision tools (e.g., a detector or depth estimator) via function calls, reason over their outputs in context, and issue high-level action primitives to a low-level controller. An MLLM agent thus places a large tool-calling language model at the center of the control loop, acting as a cognitive hub that binds perception and control through natural language.
Representative Systems.
MLLMs are playing increasingly important roles in robotics. Gemini Robotics Gemini Robotics Team (2025) integrates perception, spatial reasoning, and trajectory synthesis into a single Gemini-2.0 backbone Google DeepMind (2024), serving as an embodied brain. ManiLLM Li et al. (2024c), in a similar spirit, leverages the common-sense and reasoning capabilities of MLLMs by fine-tuning adapter modules with a chain-of-thought training paradigm, enabling accurate pose prediction and precise manipulation. These works illustrate the emerging trend of MLLMs shifting toward the role of a cognitive hub in robot systems. Hub-LLaMA Glocker et al. (2025) further builds a modular agent-orchestration system for household object management, using LLaMA 3.2 Vision Grattafiori et al. (2024) for open-vocabulary perception to ground task plans, though its limitations are not discussed.
Somewhat related to our work, Li et al. (2024a) evaluates the suitability of MLLMs as a “brain” for in-home robotics, providing a benchmark that compares models across perception, visual reasoning, and task planning. The benchmark included a few MLLMs available at the time, while newer releases were not covered—illustrating the rapid pace of progress in this area.
Strengths and Limitations.
MLLMs excel in (i) visual commonsense reasoning, leveraging extensive language priors to generalize to novel concepts beyond the reach of most specialist VLMs, and (ii) instruction following with support for fine-grained visual understanding and dynamic planning. Despite their expressive power, however, these models are (iii)  resource-intensive, posing challenges for deployment—particularly on mobile robotic platforms. We further examine the capability limits and trade-offs of MLLMs (Section 3).
3 Case Studies on Instruction Grounding
Figure 3: Experimental setup for two case studies in a cluttered tabletop environment. The top row shows egocentric video data collected for the manipulation case study. The bottom row is an example setup for the instruction grounding task, including an annotated visual prompt paired with complex instructions in three forms: implicit, explicit with attributes and spatial references.
Natural language instruction grounding translates user intent into actionable goals within a visual scene. Our case study evaluates grounding performance under challenging cross-modal disambiguation, highlighting trade-offs and offering guidance for efficient deployment.
Benchmark Dataset.
To isolate grounding ability from general vision priors, we design carefully controlled scenarios using household objects placed on a tabletop. These objects are widely represented in FM training corpora, while the tabletop setup minimizes variation in lighting and camera angles—ensuring the evaluation primarily reflects grounding performance.
We curated a new Instruction Grounding benchmark (Fig. 3) consisting of images with multiple household objects, each tagged with numbers as visual prompts and paired with instructions targeting visual commonsense and cross-modal disambiguation. For example, “pick up the red-capped marker” requires attribute reasoning to choose among markers, while “grasp the cup in front of the screwdriver” tests spatial reasoning (Appendix Table 4). These tasks highlight how grounding errors can lead to execution failures in embodied systems.
Model Easy Medium Hard Avg Specialist VLMs GroundingDINO-86M 0.518 0.357 0.349 0.408 GroundingDINO-145M 0.443 0.320 0.355 0.372 Proprietary MLLMs Gemini2.5-Pro-Exp 0.904 0.765 0.793 0.821 Gemini2.0-Flash 0.884 0.738 0.678 0.767 GPT-5-auto 0.881 0.829 0.760 0.823 GPT-5-mini 0.749 0.737 0.776 0.754 GPT-4o 0.814 0.745 0.683 0.747 GPT-4o-mini 0.803 0.722 0.604 0.710 o4-mini 0.721 0.769 0.710 0.733 Open-source MLLMs Llama-3.2V-90B 0.722 0.701 0.657 0.693 Llama-3.2V-11B 0.583 0.569 0.547 0.566 Llama-4-Maverick 0.698 0.576 0.634 0.636 Llama-4-Scout 0.776 0.615 0.624 0.672 Qwen2-VL-72B 0.686 0.614 0.558 0.619 Gemma-3-27B 0.452 0.384 0.267 0.368 DS-Janus-Pro-7B 0.444 0.330 0.317 0.364 Phi-3.5-Vision-4.2B 0.291 0.357 0.205 0.284 Base MLLMs in VLAs PaliGemma-3B ($\pi_{0}$) 0.118 0.07 0.05 0.079 QwenVL-3B (NORA) 0.519 0.535 0.577 0.543
Table 2: Object grounding performance of specialist VLMs and MLLMs across cluttered scenes of varying complexities, with macro accuracy reported.
Figure 4: Performance of complex instruction grounding across modular VLM pipeline and MLLMs. Macro accuracy is reported across instruction types—implicit, attribute-based, and relationship-based. Subfigures show (a) proprietary models and (b) open-source models along with their Int4-quantized variants.
Zero-Shot Object Grounding.
We start with a necessary first question for instruction grounding: Can FMs reliably recognize objects in cluttered, open-world scenes? Table 2 highligh key observations and point to opportunities for VLA design.
  • Despite their popularity in modular pipelines, GroundingDINO achieves around 0.3-0.4 accuracy, as it struggles with featureless objects, e.g. ‘the can’. Moreover, it is brittle in open scenes, e.g. a ‘screwdriver’ is constantly recognized as a ‘marker’, which instead is an easy case for MLLMs which embodied large volume of visual commonsense (Appendix Fig. 10).
  • Gemini 2.5-Pro and GPT-5 achieve top performance with an average score of 0.82, followed by Gemini 2.0-Flash and GPT-4.5. Open-source models continue to lag behind, with LLaMA 3.2-Vision 90B reaching 84% of the performance of top proprietay models.
  • Small- and medium-scale models (e.g., Gemma-27B, Phi-Vision) generally fall below the specialist threshold, underscoring their limitations for fine-grained object grounding in open scenes. Notable exceptions include GPT5-mini (0.754) and QwenVL-3B (0.543), which offer favorable speed–accuracy trade-offs, achieving reasonable performance without high API costs or heavy memory footprint of larger models.
  • PaliGemma-3B in $\pi_{0}$ struggles to follow structured prompts and can only follow simpler instructions. Its grounding capability is highly limited, often detecting just a single object per scene. Reported accuracy is therefore approximate, based on a looser evaluation criterion. While QwenVL-3B adopted in NORA, reaches higher grounding accuracy, which may partially explain its stronger performance on out-of-distribution manipulation tasks (Appendix Table 6).
  • Two design questions emerges: (i) how should perception modules be selected or adapted to improve instruction-following in embodied systems? and (ii) can compact, grounding-capable perception modules be distilled from large-scale models to support efficient deployment?
Zero-Shot Complex Instruction Grounding.
This task is framed as a multiple-choice problem, where the model selects the correct object index in a cluttered scene given three types of natural language instructions: implicit, attribute-based, and relationship-based. Each type probes a distinct grounding challenge (Fig. 4). As a baseline, we evaluate a modular VLM pipeline in which GPT-4 parses the instruction to infer likely targets, queries GroundingDINO to detect candidate objects, and selects from the detected boxes—essentially guessing without directly perceiving the scene.
[…]
  • Relational Reasoning Remains Challenging. This category requires resolving referential ambiguity through implicit chain-of-thought reasoning: grounding objects, modeling spatial relationships, and disambiguating targets (e.g., identifying the correct mug among many based on “next to something”). Accuracy drops significantly nearly across all models. GPT-5, Gemini 2.5-Pro and o4-mini achieve accuracy above 0.80—demonstrating the benefits from embodied training data and strong reasoning capabilities. Notably, o4-mini is a medium-sized model, yet it outperforms larger models like GPT-4o on relational instructions—suggesting that structured reasoning may help close, or even overcome the performance gap brought by different model scales.
  • Instruction-Dependent Quantization Effects. INT4 quantization reduces the model size by over 70%, making it an attractive choice for deployment. In Llama 3.2 Vision, we observe that it disproportionately impacts implicit and relational instruction grounding, indicated by the relative accuracy drop of $14\%-17\%$, while attribute grounding is more robust with only 4% loss. Despite reduced precision, quantized 11B models offer a speed–accuracy balance for low-resource settings. Our findings underscore the need for fine-grained quantization strategies that preserve the most important high-level reasoning capabilities under resource constraints.
Models LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
OpenVLA finetuned 84.7 88.4 79.2 53.7 76.5
$\boldsymbol{\pi_{0}}\$finetuned 96.8 98.8 95.8 85.2 94.15
$\pi_{0}\text{-FAST}$ finetuned 96.4 96.8 88.6 60.2 85.5
SpatialVLA finetuned-AC 88.2 89.9 78.6 55.5 78.1
NORA-finetuned 85.6 87.8 77.0 45.0 73.9
NORA-finetuned-AC 85.6 89.4 80.0 63.0 79.5
NORA-Long-finetuned 92.2 95.4 89.4 74.6 87.9
Table 3: Success rates (%) on the LIBERO Simulation Benchmark across four task suites, each evaluated over 500 trials. Results for SpatialVLA are from Qu et al. (2025); Results for $\pi_{0}$ are from Black et al. (2024), using pretrained models on LIBERO benchmarks. “AC” denotes the use of action chunking. The comparison in the Appendix highlights its impact on performance. The finetuned $\pi_{0}$ model achieves the highest performance.
4 Case Studies on Robotic Manipulation
Now we shift the focus to skill adaptation. In an ideal deployment scenario, a pretrained VLA—already endowed with broad visuomotor skills—should be retargeted to a new manipulation task with minimal data and fast convergence. We use fine-tuning, the standard practice for adaptation, as a probing lever to evaluate how the state-of-the-art VLA models adapt to new tasks and deployment conditions.
Given the scale of VLAs, we compare partial fine-tuning, which leverages our curated probing dataset (Fig. 3) and its inherent distribution bias to study convergence behavior, and full fine-tuning, which uses large-scale datasets to minimize the training loss. Our evaluation focuses on three key aspects: (i) training dynamics—how quickly and smoothly training converges; (ii) generalization—how well the resulting policies perform on various tasks; and (iii) robustness—how well the resulting policies handle environmental distractors. Our experiments highlight the performance of VLA models in different settings, offering practical suggestions for practitioners who have to adapt large VLAs under tight data, time and compute budgets.
Real-World Skill Adaptation.
Our fine-tuning process consists of two stages: (1) To assess convergence on tasks beyond the coverage of common pretraining datasets, we collected a custom dataset (Appendix A.1) focused on screwdriver-picking in cluttered tabletop scenes. This task introduces both object- and scene-level distribution gaps, as screwdrivers and dense clutter are largely absent from Open-X-Embodiment Collaboration et al. (2024). We used this dataset to partially fine-tune generalist VLA models, and to train compact task-specific models such as Diffusion Policy (DP) and Action Chunking Transformer (ACT) from scratch. (2) For full fine-tuning, we leveraged larger benchmarks—Open-X-Embodiment and the simulated LIBERO dataset Liu et al. (2023)—to fully fine-tune RT-1, OpenVLA, SpatialVLA, and NORA, and compared their performance.
Figure 5: Partial fine-tuning results for VLAs (OpenVLA and $\pi_{0}$) compared with training Diffusion Policy (DP) and ACT from scratch on our dataset. VLAs require more epochs to converge and show higher performance variance.
Figure 6: Success rates of fully fine-tuned VLAs on multi-object pick-and-place, out-of-distribution object manipulation, and spatial relationship reasoning tasks. NORA achieves the highest performance.
  • Partial Fine-tuning. We observe that DP and ACT converge stably with low training variance (Fig. 5). In contrast, generalist models such as OpenVLA and $\pi_{0}$ require far more iterations to reach comparable accuracy and exhibit greater variance. This difference reflects model scale: compact task-specific policies like DP and ACT contain only tens of millions of parameters, whereas generalist policies such as OpenVLA (7B) and $\pi_{0}$ (3B) rely on billion-scale backones, making optimization slower and less stable. Notably, while DP attains lower loss by fitting directly to noise, it still requires additional training to produce coherent action sequences even after loss convergence.
  • Full Fine-tuning. The fine-tuned VLAs are evaluated on three tasks: (1) out-of-distribution object manipulation, (2) spatial relationship reasoning, and (3) multi-object pick-and-place. In task (1), both NORA and OpenVLA succeed, whereas SpatialVLA fails due to incorrect affordance-point estimation. In task (2), NORA follows instructions correctly, while OpenVLA fails and SpatialVLA shows unstable performance. In task (3), only NORA executes the task successfully, while the other models fail to complete it reliably (Fig. 6).
Simulated Skill Adaptation.
We compare model performance on simulation benchmarks and real robot deployments (Table 3). The results show that the finetuned $\pi_{0}$ model outperforms all baselines across tasks, achieving the highest success rates on Spatial and Object tasks while maintaining strong performance on Goal and Long-horizon tasks. Furthermore, the ablation study of action chunking (AC) on NORA demonstrates that AC consistently enhances performance across most simulation tasks. Notably, a significant drop in performance is observed when transferring from simulation to the real world (Appendix Table 6).
Robustness to Perturbations.
We assess robustness by introducing distractor objects into the environment. As shown in Appendix Table 8, both OpenVLA and NORA degrade substantially under these perturbations, underscoring their sensitivity to novel conditions.
Key Takeaways.
Current VLAs still face significant limitations in the following areas:
  • Adaptation and Generalization. A generic robotic policy is expected to adapt rapidly to datasets with distributional shifts. However, our partial fine-tuning results show that, given their large model capacities and the limited size of task-specific datasets, current VLAs fail to achieve efficient adaptation. While fine-tuning improves performance, it demands extensive data and prolonged training, which is usually impractical for many real-world scenarios.
  • Robustness. Robustness to distribution shifts without finetuning remains a critical challenge. Results reveal substantial degradation when encountering unseen objects and during sim-to-real transfer, underscoring the fragility of current VLA models in dynamic and unpredictable environments.
These findings indicate that although VLAs hold significant promise, they remain constrained by poor data efficiency, slow adaptation, and brittle robustness. Bridging these gaps will require both algorithmic advances—such as more parameter-efficient adaptation, bias-resistant pretraining, and stronger perception backbones—and system-level improvements in data collection and augmentation. Addressing these limitations is crucial if VLAs are to evolve from research prototypes into reliable, deployable policies for real-world robots operating under uncertainty.
5 Constraints and Future Directions
Despite the promise of foundation models (FMs) for enabling embodied agents to perform daily tasks, several critical constraints still hinder their reliable deployment:
Data Scarcity.
Unlike natural language data that abundantly available on the internet, robotic datasets are costly to collect due to hardware wear, safety risks, and labor-intensive demonstrations. A key direction is improving data efficiency through parameter-efficient adaptation, imitation from unlabeled interaction, and self-supervised pretraining. Complementary approaches include leveraging high-fidelity simulators and developing robust sim-to-real transfer pipelines to reduce reliance on large real-world collections.
Efficient Inference.
VLAs place heavy computational burdens on robotic platforms, leading to bottlenecks in inference speed. The constraints motivate research on lightweight architectures and efficient decoding strategies that can sustain performance while satisfying the real-time requirements of embodied control.
Explainability and Safety.
Most FMs lack explicit mechanisms for interpretability or safety guarantees—factors that are crucial for deployment in high-stakes or unstructured environments. These models may output confident but incorrect actions when faced with out-of-distribution inputs or adversarial perturbations. Moreover, without built-in constraints to enforce ethical and operational boundaries, VLAs risk misinterpreting ambiguous instructions in ways that compromise human intent or physical safety.
Limitations
For VLA generalization study, this work focuses primarily on task-level performance and does not extensively examine generalization across diverse robot morphologies. Yet this remains a critical challenge for real-world deployment: robots with different embodiments—such as bimanual manipulators, humanoid robots, or mobile platforms—require distinct control protocols and safety constraints. The absence of a generic policy that adapts seamlessly across such morphologies limits the universality of current VLA approaches.
This work also does not directly investigate the grounding of instructions that are open-ended or ambiguous. Existing VLAs are trained largely on curated datasets that map well-structured commands to specific actions, but this reliance constrains their semantic understanding. Consequently, when faced with vague or out-of-distribution instructions, they often fail to infer reasonable behaviors. Addressing this limitation will require integrating richer language-understanding modules and more diverse training data to improve robustness in handling underspecified or ambiguous input.
Acknowledgment
This work is supported by the National Robotics Programme (Award #M23NKBK0091, #M25N4N2009). This work is also supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-005), NTU SUG project #025628-00001:Post-training to Improve Embodied AI Agents, and A*STAR SERC CRF funding to Cheston Tan. We thank Wei Qi Toh and Ray Prithviraj for their assistance with robot experiments, and Shaohua Li for manuscript polishing.
Appendix
Appendix A Benchmark Dataset
A.1 Cluttered Tabletop Manipulation Dataset
To evaluate the finetuning behavior of various VLA models under distribution shift, we constructed a custom cluttered tabletop environment using a UR5 robotic arm with a wrist-mounted RealSense RGB-D camera. This setup differs from all existing configurations in the Open-X-Embodiment dataset. Demonstrations for a screwdriver-picking task—amid distractor objects—were collected via teleoperation using a SpaceMouse device. In total, we gathered 163 demonstration episodes [^3] https://huggingface.co/datasets/bittdy/pick_screw. Each episode began with a randomized initial robot pose, followed by an attempt to grasp the target screwdriver.
A.2 Complex Instruction Grounding Dataset
[…]
Words Positional:       left, right, between, beside, near, far, front, behind Directional:     aligned with, perpendicular to Instruction hand over the screwdriver [on the left of] the red ball. pass me the screwdriver [aligned with] the marker.
Table 4: Template words and corresponding examples of generated relation-based instructions for case studies.
For the complex instruction grounding task, including an annotated visual prompt paired with complex instructions in three forms: implicit, explicit with attributes and spatial references. Additionally, the dataset includes multi-turn questions that refer to more than one object, enabling foundation models to ask clarifying questions to identify the correct object.
  • Implicit Instructions. Here, objects are not explicitly mentioned by name or attributes but are instead described by their functions. This category evaluates the VLMs’ ability to infer the correct object based on its use. For example, the dataset includes instructions referring to objects like scissors, screwdrivers, and rulers based on their respective functions.
    Figure 7: Example of visual prompts
  • Explicit Attributes. In this category, instructions prompt VLMs to identify objects belonging to a category with multiple instances, where each instance can be uniquely identified by explicitly mentioned attributes. In Fig. 7, the beige mug and the gray mug are included because they are unique when described with attributes. However, objects like the black mug or scissors are excluded. This is because there are two identical black mugs, making them non-unique, and there is only one pair of scissors, which does not require attributes for identification.
  • Explicit Relationships. In this category, instructions describe objects by their spatial relationships to other objects in the image. We ensure that each referenced object is unique within the image. For example, the measuring cup to the right of the screwdriver uniquely identifies the object. These instructions are designed to test the VLMs’ ability to comprehend and resolve location-based relationships.
  • Multi-Referent Instructions. This category includes instructions that correspond to multiple valid objects in the scene. For example, in Fig. 7, an instruction like “give me a mug” may refer to several similar items. In such cases, we annotate the data with all candidate object indices, e.g., [2, 7, 18], indicating the set of plausible referents.
Figure 8: Dataset breakdown by Instruction Types.
Model Easy Medium Hard
im attr rel im attr rel im attr rel
VLM-LLM 0.050 0.516 0.131 0.010 0.336 0.186 0.000 0.318 0.174
Gemini-2.5-Pro 0.778 0.889 0.830 0.847 0.814 0.815 0.985 0.784 0.858
Gemini-2.0 0.833 0.889 0.774 0.819 0.721 0.642 1.000 0.668 0.469
GPT-5-auto 0.950 1.000 0.867 0.903 0.960 0.899 0.917 0.916 0.816
GPT-5-mini 0.850 1.000 0.894 0.819 0.956 0.823 0.958 0.759 0.813
GPT-4.5 0.944 0.889 0.894 0.917 0.838 0.722 0.958 0.719 0.698
GPT-4o 0.850 1.000 0.778 0.819 0.948 0.680 0.901 0.697 0.469
o4 0.950 1.000 0.907 0.847 0.988 0.837 0.917 0.809 0.804
4o-mini 0.750 0.717 0.550 0.764 0.771 0.596 0.750 0.382 0.248
GPT-4V 0.650 0.750 0.598 0.750 0.737 0.662 0.625 0.417 0.455
Qwen2-VL 0.800 0.917 0.830 0.792 0.756 0.738 0.875 0.700 0.529
LLaMA 3.2 Vision 90B 0.750 0.850 0.704 0.708 0.853 0.711 0.875 0.491 0.521
LLaMA 3.2 Vision 90B-Q4 0.800 0.667 0.598 0.625 0.719 0.554 0.542 0.464 0.300
LLaMA 3.2 Vision 11B 0.650 0.667 0.631 0.764 0.710 0.556 0.833 0.536 0.342
LLaMA 3.2 Vision 11B-Q4 0.650 0.567 0.502 0.694 0.757 0.555 0.542 0.498 0.450
Table 5: Performance on the complex instruction grounding task. Abbreviations: im denotes implicit instruction, attr denotes attribute-based instruction, and rel denotes relation-based instruction.
A human-in-the-loop process was employed to ensure high-quality data collection.
  • Initial Object Identification: We used GPT-4o to identify objects in an image and referring them by type, explicit attributes, and detailed location relations.
  • Human Verification. The authors of this paper reviewed and modified the outputs to ensure their correctness.
  • Instruction Generation. After verification, GPT-4 was tasked with generating simple, clear instructions for different objects.
  • Final Review. These instructions underwent another round of verification to ensure clarity and accuracy.
This high-quality dataset consisting of 473 instructions, with a detailed breakdown of each instruction type presented in Fig. 8.
Figure 9: Examples of Instruction Grounding. (a) “the marker on the left”, (b) “the marker aligned with the ruler”.
Figure 10: Examples of Object Grounding. (a) “ball”, (b) “screwdriver”, (c) “marker pens”, (d) “blue ball”.
Appendix B Grounding Experiments
B.1 Complex Instruction Grounding for Goal Specification
Cross-modal Disambiguation represents a particularly challenging component of goal specification. To quantify the model capability in this dimension, we employed attribute-based and relative relationship instructions to uniquely identify a target among multiple candidates. The goal specification task is formulated as follows:
Given a visual input $I\in\mathbb{R}^{H\times W\times 3}$ and an instruction $t$, the objective is to predict the target object according to $o^{*}=\arg\max_{o\in\mathcal{O}}P(o\,|\,I,t;\theta_{FM})$, where $o^{\star}\in\mathcal{O}$ and $\mathcal{O}$ denotes the set of candidate objects. Models are evaluated using macro-average accuracy metric.
B.2 Failure Cases of Specialist VLM Pipeline
Grounding DINO, despite popular for zero-shot detection, is not robust in open scenes. It successfully detected “blue ball” while failed to detect “ball”, indicating its reliance on visual features. Similarly, featureless metal cans pose a great challenge for Grounding DINO, which were almost omitted in the detection results.
For complex instruction grounding, Grounding DINO and GPT-4 were chained together to “guess” the target by the LLM based on the candidate bounding boxes. The failure cases were illustrated in the Fig. 9 and Fig. 10.
B.3 Multimodal LLMs Performance
The performance of Multimodal LLMs on complex grounding across Easy, Medium and Hard groups are shown in Table 5.
Appendix C Manipulation Experiments
OpenVLA and $\pi_{0}$ were partially fine-tuned on this dataset, while Diffusion Policy (DP) and Action Chunking Transformer (ACT) were trained from scratch. Due to the limited size of our custom dataset, full fine-tuning of RT-1, OpenVLA, SpatialVLA, and NORA was performed using the Open-X-embodiment and LIBERO datasets. We evaluated model performance on both a real-world WidowX robotic platform and the LIBERO simulation benchmark.
C.1 Fine-tuning details for VLA
Partial fine-tuning was conducted on a single NVIDIA A6000 GPU (48 GB VRAM) over a period of three days. To ensure a fair comparison, a batch size of 1 was used across all models. The results are presented in Fig. 5.
[…]
Figure 11: Real-world robot environments and task setups. We evaluate these models across 9 diverse tasks to assess its instruction understanding, spatial reasoning, and multi-task motion planning capabilities.
Category Task RT-1 OpenVLA SpatialVLA NORA
Multiple objects Put the red bottle and the hamburger in the pan 0 20 0 40
Put the carrot and hotdog in pot 0 0 0 30
Put the corn and carrot in the pan 0 30 0 30
OOD object put carrot in pot 0 80 20 90
Put banana in pot 1 40 0 90
Put the blue cube on the plate 0 50 0 70
Spatial Put the pink toy at the right corner 0 60 30 60
Put the blue cube on the right plate 0 30 0 20
Move the banana close to the pan 30 50 50 80
Average 4.4 40 11.1 56.7
Table 6: Task performance comparison across different categories and models.
C.2 Impact of Action Chunking
C.2.1 Action Chunking Performs on WidowX.
To investigate the effectiveness of action chunking, we selected NORA-LONG and SpatialVLA for evaluation. Tasks were chosen from three categories: (1) “put the carrot in the pot,” (2) “put the red bottle and hamburger in the pot,” and (3) “put the pink toy at the right corner.” In initial experiments, all predicted actions (5 actions for NORA-LONG, 4 actions for SpatialVLA) were executed sequentially without replanning. This frequently caused the WidowX robot to crash into the environment due to the accumulation of overly large movements.
Subsequently, we modified the execution policy to only perform the first action in each predicted chunk. This adjustment resolved the collision issue but the model performance is still degraded.
C.2.2 Action chunking improves performance in simulation.
We hypothesize that action chunking is more effective at higher control frequencies. For example, Diffusion Policy generates commands at 10 Hz, which are then interpolated to 125 Hz for execution. Similarly, OpenVLA-OFT+ employs action chunking and shows improved performance in real-world ALOHA tasks, which run at 25 Hz.
Since our real robotic platforms do not support high-frequency control, we tested this hypothesis in the LIBERO simulation environment (20 Hz). We fine-tuned both NORA and NORA-LONG on this benchmark with an action chunk size of 5, producing two variants: NORA-finetuned-AC and NORA-Long-finetuned.
Results show that NORA-finetuned-AC significantly outperforms NORA-finetuned across all LIBERO benchmarks, with a higher average success rate. Notably, NORA-Long-finetuned outperforms all baseline models (see Table 3), highlighting the benefits of pretraining with action chunking and its transferability to long-horizon tasks. However, it is important to note that LIBERO is a simulation environment and may not reflect real-world performance at high control frequencies.
Figure 12: Comparison of tasks with and without distraction.
TASK OpenVLA NORA
without Distraction
put carrot in pot 8 9
put banana in pot 4 9
put the blue cube on the plate 5 7
with Distraction
put carrot in pot 6 8
put banana in pot 6 4
put the blue cube on the plate 3 5
Table 7: Comparison of task performance between OpenVLA and NORA under conditions with and without distraction. Each value denotes the number of successful executions out of 10 trials.
C.3 Robustness to Disturbance
To evaluate robustness, we selected three straightforward tasks (shown in Fig. 12) and introduced distractor objects into the environment. Initially, both OpenVLA and NORA performed well. However, their success rates declined significantly with the introduction of distractions. This highlights the sensitivity of current VLA models to out-of-distribution disturbances. The average success rates across the three tasks are presented in Table 8, while the detailed number of successful executions out of 10 trials is summarized in Table 7.
Table 8: Average Success Rate (%) without (w/o) and with (w/) Distractors
Model w/o Distractors w/ Distractors
OpenVLA 56.7 50
NORA 83.3 56.7
Appendix D Modular Claw Machine Prototype
To facilitate the evaluation of different VLMs in robotic manipulation, we developed a voice-controlled testbed using a UR5 robotic arm [^5] https://github.com/HRItdy/claw_machine. The system architecture, shown in Fig. 13, comprises the following five modules:
  • Speech Transcription: Powered by Microsoft Azure’s speech recognition service.
  • Task Decomposition: Based on GPT-3.5 and GPT-4 using prompting paradigms adapted from ChatGPT for Robotics.
  • Object Detection: Utilizes GroundingDINO and OWL-ViT for object detection.
  • Object Segmentation: Employs Segment Anything Model (SAM) and FastSAM for segmenting detected objects.
    Figure 13: The system architecture of the testbed for VLMs.
  • Manipulation: Low-level actions are generated by GraspAnything or GraspNet.
This modular testbed enables rapid integration and benchmarking of different models within a real robotic system.